# Breaking the Curse of Multilinguality with Cross-lingual Expert Language Models

### v1

@@ 5.3 Perplexity Evaluation
[…]
@@
Table 1: Per-language and average perplexity results for the $k=8$ x-elm experiments (original XGLM and $k=1$ dense model included for comparison). Lower numbers are better. The best setting for each language is bolded per compute budget. $*{}^{*}$-IDF ensemble uses more parameters for inference than other evaluations; see Table 8 for the effect of sparsifying these ensembles on perplexity.
6 Language Modeling Experiments
We now test the effectiveness of sparse language modeling in the multilingual setting. First, we determine the optimal number of clusters for our given compute budget and dataset (§6.1). We then demonstrate that x-elms outperform comparable dense models on seen languages (§6.2) and more effectively adapt to new, unseen languages (§6.3). Finally, we examine the effect of sparse training on forgetting previously-held knowledge of languages in specific x-elm experts (§6.4).
6.1 Choosing the Number of x-elms
We first consider which choice of $k$ clusters gives the best multilingual language modeling performance. Figure 4 compares the choice of $k=1,4,8,16$ x-elms when trained on 10.5B tokens. [^7] The $k=16$ setting is equivalent to training monolingual experts for every language. Full results are in Table 1 for $k=8$ and Appendix C for $k=4$ and $k=16$. $k=8$ is the best-performing setting on 75% of languages when clustering with TF-IDF and for 15 of the 16 pretraining languages when clustering by language similarity. Furthermore, typological clustering consistently outperforms TF-IDF.
These experiments indicate that, for the budget we evaluate, the best overall x-elm setting is bilingual models (k=8) clustered by language similarity. This result is surprising, as it is intuitive to assume that simply continuing to pretrain each expert on a single language (i.e., the $k=16$ setting) would lead to better perplexity. We find that one language, Swahili, does benefit from the monolingual $k=16$ setting—possibly because Swahili is paired with a distant language (Vietnamese) by the typological clustering process. However, perplexity is higher in the $k=16$ setting for all other languages, and in some cases even underperforms the dense ($k=1$) model.
@@ 6.2 Perplexity Results on Seen Languages
Figure 5: Comparison of PPL improvements per language over XGLM-1.7B (circle) and dense baseline (triangle) against the training data quantity (for typologically clustered experts).
@@ 6.3 Unseen Languages and Modeling New Languages with x-elm
We also examine how well x-elm performs on held-out languages as well as adapts to new languages. Specifically, we consider both zero-shot evaluation and further training of x-elm on four languages not included in the original XGLM seed model: Azerbaijani (AZ), Hebrew (HE), Polish (PL), and Swedish (SV). [^8] Data for these languages is also obtained from mC4, with the same preprocessing as other languages in our experiments.
@@
Table 2 presents the unseen target language perplexities in the XGLM and x-elm Training columns. We find that the original XGLM model performs poorly on the new languages, particularly those less related to XGLM’s highest-resourced ones (i.e., AZ and HE). While these perplexities remain high in the dense model and TF-IDF ensembles, training (on other languages) with x-BTM provides some performance improvements over the seed model.
@@
Adapting x-elm to new languages
We now consider how well Hierarchical Multi-Round training (HMR) works for language adaptive pretraining (LAPT, Chau et al., 2020), which incorporates new target languages into the continued pretraining process. Here, we group each target language with a higher-resource donor language already in our pretraining set; these are assigned with the language similarity metric used for typological clustering. We seed each new language’s expert with an expert specialized to that language’s donor; the new expert is then trained on the donor/target language pair. For HMR inference, we evaluate perplexity with the expert trained on that target language. [^9] We also evaluate the donor languages to see what benefit, if any, they receive from the adaptation process.
All of the LAPT settings provide considerable improvements on the new target languages over the unseen language experiments (Table 2, LAPT columns). The HMR setting outperforms continued dense training on every new language. Furthermore, HMR training removes the risk of catastrophic forgetting Yogatama et al. (2019) in other LAPT schemes, as this process adds new experts to x-elm rather than changing existing ones.
Figure 6: Heatmap comparing individual x-elm perplexities to the seed LM with TF-IDF (left) and Typ. (right) clustering. Positive scores indicate that the expert forgot that language. For Typ. clusters, languages that the model was explicitly trained on are grayed out.
@@
Typ. (TRG) 42.15 ${}^{\dagger}$14.354.62 ${}^{\dagger}$71.455.39 ${}^{\dagger}$28.6%
Typ. (EN) 42.43 ${}^{\dagger}$7.155.54 ${}^{\dagger}$28.655.13 14.3%
TF-IDF (Top-1) 42.55 21.4% 55.03 ${}^{\dagger}$14.355.50 ${}^{\dagger}$42.9%
@@
Table 3: Average performance and the percentage of languages where this setting outperforms the others (Win Rate) on the overlap of task evaluation languages and the x-elm target languages. The few-shot setting provides $k$=8 English demonstrations to the model and averages performance across five runs. ${}^{\dagger}$indicates (best) performance ties between two evaluation settings on a language.
6.4 x-elm Forgetting
The preceding sections evaluate x-elms as an ensemble of models by dynamically choosing the best expert for a given evaluation setting or ensembling the experts’ outputs. However, each expert is initialized with a model trained on all the languages we consider. This prompts the question: how much do individual experts forget [^10] We consider an expert to have forgetten information about a language if its perplexity on that language increases. about the languages they are not specialized to?
Forgetting occurs as x-elms become more specialized.
We compare the perplexity of each expert model on all pretraining languages to that of the seed model, XGLM-1.7B (Figure 6 for $k=8$; other settings given in Appendix C). Across the considered values of $k$, we see less forgetting in the x-elms trained on TF-IDF clusters than in those clustered typologically. For the $k=8$ expert setting, the TF-IDF experts only forget on 47.7% of settings, and when forgetting occurs, the perplexity increase over the baseline is 3.10 on average. For typologically clustered experts, these measures are 83.6% and 3.14, respectively; we observe similar trends for the $k=4$ and $k=16$ x-elms. This implies that though in some cases only small quantities of data are shared across TF-IDF clusters, these data mitigate forgetting over the hard cluster assignments made by typological clustering.
x-elms are more likely to forget certain languages.
@@ 7 In-Context Learning Experiments
We also measure whether the perplexity improvements from x-elms correspond to better performance on downstream tasks. We test the performance of our x-elms on three tasks through an in-context learning (ICL) framework, showing that the x-elm language modeling gains do translate to ICL improvements over the baseline models.
@@ 7.1 Experimental Setup
XNLI Conneau et al. (2018) is a multilingual natural language inference benchmark covering 14 of our 16 pretraining languages (excluding JA and KO). Since there are no gold training examples for XNLI, we use the test set for evaulation and sample demonstrations from the validation set.
@@
PAWS-X Yang et al. (2019) is a binary classification task that requires the model to determine whether a pair of sentences are paraphrases. This benchmark covers seven of our pretraining languages, including two (JA and KO) not covered by the other ICL benchmarks.
We compare the performance of x-elm against dense baselines in both zero- and few-shot learning settings. For all benchmarks, we evaluate on 1,000 random examples and perform five runs on different demonstration sets for few-shot settings. Unless otherwise stated, we evaluate performance on the development set and sample demonstrations from the training set; further details about the ICL evaluation protocol are given in Appendix A.2.
@@ 7.2 Results
We evaluate our best x-elm setting by perplexity—$k$=8 experts trained on the larger compute budget of 21B training tokens—on the downstream tasks. Table 3 summarizes the results of these evaluations on the languages covered by the x-elm models; individual language results are given in Appendix C. The x-elm models outperform both the seed model and the compute-matched dense baseline across the three tasks and in both the zero- and few-shot evaluation settings.
Furthermore, though x-elm improves over the seed model, the dense model underperforms XGLM. This may be due to using different data from the original XGLM pretraining; data quality issues have been previously documented for mC4 Kreutzer et al. (2022); Chung et al. (2023). We also note that XNLI and XStoryCloze few-shot performance is consistently lower than in the zero-shot setting; this is a recurring issue in multilingual ICL also observed in the base model Lin et al. (2022).
@@ 8 Related Work
8.1 Multilingual Pretraining
Many variations and improvements on dense multilingual pretraining have been proposed since the introduction of multilingual BERT Devlin et al. (2019): by changing the architecture and scaling the model size up Goyal et al. (2021); Lin et al. (2022), combining additional objectives to the main LM objective Conneau and Lample (2019); Chi et al. (2022); Reid and Artetxe (2022), careful language and data curation Scao et al. (2022); Ogunremi et al. (2023), and scaling and balancing the vocabulary across the different languages Liang et al. (2023). Most relevant to our work is Pfeiffer et al. (2022), which proposes a new modular model architecture, X-MOD, that contains language-specific modules. However, many of the limitations of dense modeling persist in this architecture since the model and modules are jointly trained.
An issue common to most methods for multilingual pretraining is the curse of multilinguality Conneau et al. (2020). Wu and Dredze (2020) demonstrate that multilingual training leads to lower performance on low-resource languages than higher-resourced ones. Blevins et al. (2022) find that multilingual models forget information previously learned during training, which they postulate is due to this phenomenon; Wang et al. (2020) similarly suggest that this effect occurs due to training dynamics. More recently, Chang et al. (2023) presented a controlled study of the factors causing this curse that corroborates limited model capacity as the underlying cause. A primary motivation of this work is to limit the effect of this curse while maintaining the other benefits of multilingual modeling.
8.2 Adapting Multilingual Models
Another common thread of multilingual modeling research focuses on adapting an existing model to new languages. Initially, these methods continued pretraining these models with the new languages incorporated into the training regime, such as language-adaptive pretraining (LAPT; Chau et al., 2020). Other work proposed the use of adapters to update the model to new languages Pfeiffer et al. (2020); notably, Faisal and Anastasopoulos (2022) used similar linguistic motivations to our typological clustering to group languages into adapters. However, follow-up work found that continued pretraining outperformed adapter methods for new language adaptation Ebrahimi and Kann (2021).
@@ 8.3 Sparse Models for NLP
Sparsely activated language models (Evci et al., 2020; Mostafa and Wang, 2019; Dettmers and Zettlemoyer, 2019) route inputs through a subset of the total model parameters. Our work builds most directly on the Branch-Train-Merge (Li et al., 2022; Gururangan et al., 2023) algorithm, which results in full-model experts trained to specialize on domains of data defined by metadata or a learned clustering. This design expands both on the independent feed-forward network experts found in early Mixture-of-Experts (MoE) models (Jacobs et al., 1991) and on DEMix layers (Gururangan et al., 2022), which routes sequences to per-layer feed-forward experts based on metadata.
Other MoE models have recently been applied to multilingual settings. Pfeiffer et al. (2022) develop a multilingual expert model with language-specific routing, and Kudugunta et al. (2021) develop a machine translation model with routing determined by the source-target language pair or the target language. Similarly to BTM, Jang et al. (2023) trains experts specialized to different tasks, including five machine-translation language pairs, which can be merged with other task experts.
@@ 9 Conclusion
This work presents an approach to mitigate the curse of multilinguality by extending sparse language modeling to the multilingual setting with x-elms (cross-lingual expert language models). We find that x-elms achieve better perplexity over standard, dense language models trained with the same compute budget; these experts can also be efficiently adapted to new languages without the risk of catastrophic forgetting. x-elms also present other benefits over dense models for multilingual modeling, such as not disproportionally benefitting high-resource languages over lower-resourced ones. Finally, we show that these language modeling improvements transfer to downstream tasks.
While our experiments show that x-elm outperforms dense LMs, we foresee many avenues of future work to further tailor sparse modeling to multilinguality. These include better methods for data allocation—such as clustering methods that leverage cross-lingual signal— and algorithmic improvements to better allocate compute and more effectively ensemble models at inference. By proving the efficacy of sparse language modeling in the multilingual setting, we hope to inspire future work in this vein that fairly models every language while leveraging the potential of cross-lingual learning.
@@
One limitation of this work is that we focus on examining the effect of training x-elms with x-BTM in a limited number of settings, training languages, and data sources fixed; this is due to limited computational resources. Therefore, the proposed methods should be verified in other settings. In particular, we hope to examine how x-elm performs at scale when using larger experts, more languages, and larger training budgets.
@@ Appendix A Additional Experimental Details
# Tokens k # GPUs # updates grad acc.
10.5 B 1 8 20,000 32
4 4 20,000 16
8 4 20,000 8
16 2 20,000 8
21.0 B 1 8 40,000 32
4 4 40,000 16
8 4 40,000 8
16 2 40,000 8
Table 4: Overview of the total compute budget and resources used for different x-elm experiments. k is the number of experts, # GPUs indicates the number of GPUs used to train each expert, and grad acc. gives the number of gradient accumulation steps used.
Language ${}^{\dagger}$ Size (%) XGLM Size
@@ 21.0 B 1 8 40,000 32
Table 5: The frequencies and relative percentages of different languages in our training corpus (${}^{\dagger}$a subsampled version of mC4) and in the XGLM pretraining corpus, CC100-XL (as reported in Lin et al. (2022)). Sizes of data are reported in gigabytes (GiB). EN, ES, FR, and RU are downsampled to the first 1,024 mc4 shards for those languages.
A.1 Pretraining
@@
A.2 In-Context Learning
We reimplement the evaluation protocol from Lin et al. (2022), where the model scores multiple versions of every example (with the different possible labels filled in), and the label of the highest-scoring version is considered as the model’s prediction. We use the English prompt formats and evaluation protocols developed for the seed LM of our experts, XGLM, for the downstream tasks of XNLI, XStoryCloze, and PAWS-X. The prompt templates we use are reproduced in Table 6.
In the few-shot experiments, we perform five evaluation runs with different demonstration samples and reported the average performance. All few-shot experiments are performed with eight random demonstrations. As we are testing the cross-lingual abilities of x-elm, these demonstrations are in English for every target language.
Lang. k=4 Experts k=16 Experts k=8
[…]
AR 16.85 15.29 14.99 15.03 15.00 15.60 15.67 15.40 14.36
BG 11.31 10.44 10.39 10.39 10.42 11.10 10.70 10.31 10.18
DE 15.53 14.02 13.85 13.89 13.71 14.71 14.43 14.5 12.13
EL 10.44 9.40 9.36 9.33 9.28 9.72 9.64 9.41 9.05
EN 14.37 12.88 12.64 12.71 12.78 13.60 13.23 13.27 12.60
ES 16.02 14.13 13.93 13.96 14.06 14.83 14.58 14.59 13.75
FR 13.12 11.78 11.62 11.65 11.55 12.38 12.13 12.15 11.05
HI 18.28 14.28 14.22 14.21 12.64 16.11 15.67 13.86 10.98
JA 14.57 12.31 12.23 12.12 11.73 13.39 13.14 13.18 11.01
KO 8.82 7.78 7.81 7.77 7.70 8.14 8.09 7.75 7.56
RU 13.43 12.52 12.30 12.33 12.46 12.96 12.76 12.82 11.95
SW 19.85 18.70 18.61 18.62 18.19 19.38 19.13 16.43 18.30
TR 17.81 15.34 14.85 14.96 14.81 15.67 15.78 15.52 13.47
UR 14.38 13.45 13.56 13.73 13.18 13.88 13.87 12.65 12.46
VI 13.07 11.39 11.43 11.21 10.32 11.85 11.65 11.59 10.21
ZH 17.91 13.74 13.38 13.70 13.11 14.65 14.95 13.58 11.66
Avg. 14.74 12.97 12.82 12.85 12.56 13.62 13.46 12.94 11.98
Table 7: Per-language and average perplexity results for the $k=4$ and $k=16$ x-elm experiments (original XGLM and $k=1$ dense model included for comparison). Lower numbers are better. Each x-elm setting is trained on 10.5B tokens. $*{}^{*}$-IDF ensemble uses more parameters for inference than other evaluations. ${}^{\dagger}$he HMR models are initialized from an existing expert and trained for 10.5B more tokens.

### v2

@@ 6.1 Choosing the Number of x-elms
XGLM Dense TF-IDFtop1 TF-IDFens Typ. Dense TF-IDFtop1 TF-IDFens Typ.
@@
Table 1: Per-language and average perplexity results for the $k=8$ x-elm experiments (original XGLM and $k=1$ dense model included for comparison). Lower numbers are better. The best setting for each language is bolded per compute budget. TF-IDF ensemble uses more parameters for inference than other evaluations; see Table 8 for the effect of sparsifying these ensembles on perplexity.
@@ 6.2 Perplexity Results on Seen Languages
Figure 4: PPL improvements per language over XGLM-1.7B (circle) and dense baseline (triangle) against the training data quantity (for typ. clustered experts).
@@ 6.3 Unseen Languages and Modeling New Languages with x-elm
A common method for applying multilingual language models to new settings is language-adaptive pretraining (LAPT; Chau et al., 2020)), as this reuses multilingual knowledge in existing LMs while training on the target language(s). We now examine how well x-elm performs in this paradigm by (1) evaluating x-elm perplexity on unseen languages and (2) adapting an existing x-elm set to new languages. Specifically, we consider both zero-shot evaluation and further training of x-elm on four languages not included in the original XGLM seed model: Azerbaijani (AZ), Hebrew (HE), Polish (PL), and Swedish (SV). [^9] Data for these languages is also obtained from mC4, with the same preprocessing as other languages in our experiments.
@@
Table 2 presents the unseen target language perplexities in the XGLM and x-elm Training columns. We find that the original XGLM model performs poorly on the new languages, particularly those less related to XGLM’s highest-resourced ones (i.e., AZ and HE). While the perplexities remain high in the dense model and TF-IDF ensembles, continued training on other languages improves the seed model.
@@
Typ. (TRG) 42.15 14.3% 54.62 71.4% 55.39 28.6%
Typ. (EN) 42.43 7.1% 55.54 28.6% 55.13 14.3%
TF-IDF (Top-1) 42.55 21.4% 55.03 14.3% 55.50 42.9%
@@
Table 3: Average performance (Acc.) and the percentage of languages where this setting outperforms the others (Win Rate) on the overlap of task evaluation languages and the x-elm target languages. The few-shot setting provides $k$=8 English demonstrations to the model and averages performance across five runs. Indicates (best) performance ties between two evaluation settings on a language when calculating the win rate.
Adapting x-elm to new languages
We now consider how well Hierarchical Multi-Round training (HMR, §4) works for language adaptive pretraining (LAPT, Chau et al., 2020), which incorporates new target languages into the continued pretraining process. Here, we group each target language with a higher-resource donor language already in our pretraining set; these are assigned with the language similarity metric used for typological clustering. We seed each new language’s expert with an expert specialized to that language’s donor; the new expert is then trained on the donor/target language pair. For HMR inference, we evaluate perplexity with the expert trained for that target language; we also evaluate on the donor languages to see what benefit, if any, they receive from the adaptation process.
All of the LAPT settings provide considerable improvements on the new target languages over the unseen language experiments (Table 2, LAPT columns). The HMR setting outperforms continued dense training on every new language. Furthermore, HMR training removes the risk of catastrophic forgetting Yogatama et al. (2019) in other LAPT schemes, as this process adds new experts to x-elm rather than changing existing ones. [^10] This forgetting of known languages occurs in our dense LAPT baseline, with perplexity decreasing by 1.91 points on average for languages not included in the adaptation setting.
@@ 7 In-Context Learning Experiments
We also measure whether the perplexity improvements in x-elms correspond to better performance on downstream tasks. We test x-elms on three tasks in an in-context learning (ICL) framework, showing x-elm language modeling gains do translate to ICL improvements over the baseline models.
@@ 7.1 Experimental Setup
XNLI Conneau et al. (2018) is a multilingual natural language inference benchmark covering 14 of our 16 pretraining languages (excluding JA and KO). Since there are no gold training examples for XNLI, we use the test set for evaluation and sample demonstrations from the validation set.
@@
PAWS-X Yang et al. (2019) is a binary classification task that requires the model to determine whether a pair of sentences are paraphrases. This benchmark covers seven of our pretraining languages, including two (JA and KO) that are not covered by the other ICL benchmarks.
We compare x-elms against dense baselines in zero- and few-shot settings. For all benchmarks, we evaluate on 1,000 random examples and perform five runs on different demonstrations for few-shot learning, using English demonstrations for every language to test cross-lingual transfer. Further details about the ICL evaluation are in Appendix A.2.
@@ 7.2 Results
We evaluate our best x-elms by perplexity—$k$=8 experts trained on the larger compute budget of 21B training tokens—on their in-context learning abilities (Table 3). [^11] Full results for each task are given in Appendix Tables 9, 10, and 11. Here, we consider two metrics to summarize model performance across languages: Acc. is the average accuracy of each model for that ICL task across evaluation languages, and Win Rate is the percentage of languages where that model achieves the highest score out of the considered models; if two models get the highest score, they are both considered to win that setting.
We find that x-elms outperform both the seed and compute-matched dense baselines across the three tasks and in both the zero- and few-shot evaluation settings. Furthermore, though x-elm improves over the seed model, the dense model underperforms XGLM. This may be due to using different data from the original XGLM pretraining, as data quality issues have been documented for mC4 Kreutzer et al. (2022); Chung et al. (2023). We also note that few-shot ICL performance on XNLI and XStoryCloze is consistently lower than in the zero-shot setting; this is a recurring issue in multilingual ICL also observed in the seed model Lin et al. (2022).
@@ 8 Related Work
8.1 Multilingual Pretraining and Adapation
Many variants of dense multilingual pretraining have been proposed since multilingual BERT Devlin et al. (2019): changing the architecture and scaling the model size up Goyal et al. (2021); Lin et al. (2022), adding additional cross-lingual objectives Conneau and Lample (2019); Chi et al. (2022); Reid and Artetxe (2022), and careful language and data curation Scao et al. (2022). Most similar to our work is Pfeiffer et al. (2022), which proposes an architecture, X-MOD, with language-specific modules. However, many slimitations of dense modeling persist here as the model and language modules are jointly trained.
Across most multilingual pretraining methods is the curse of multilinguality Conneau et al. (2020), particularly for lower-resource languages in massively multilingual training Wu and Dredze (2020). Blevins et al. (2022) find that multilingual models forget information previously learned during training, which they hypothesize is due to this phenomenon; Wang et al. (2020) similarly suggest that this effect is due to training dynamics. More recently, Chang et al. (2023) presented a controlled study corroborating limited model capacity as a cause of this curse. A primary motivation of our work is to mitigate this curse while maintaining the other benefits of multilingual modeling.
However, not all multilinguality is harmful to language modeling. Chang et al. (2023) show that seeing linguistically similar languages can benefit low-resource language performance; this corroborates our finding that x-elms trained on related languages outperfrom monolingual experts. In this vein, a recent direction in multilinguality has been targeted multilingual modeling, where models are trained on data from the same language family Ogueji et al. (2021); Ogunremi et al. (2023); Ljubešić et al. (2024); Downey et al. (2024).
We also consider how x-elm can be used for language adaption. The most common method, language-adaptive pretraining (LAPT; Chau et al., 2020), continues multilingual pretraining with new languages incorporated into the training regime. Other work proposed using adapters to update the model with new languages Pfeiffer et al. (2020); notably, Faisal and Anastasopoulos (2022) used similar linguistic motivations to our typological clustering to group languages for adapters. However, Ebrahimi and Kann (2021) found that LAPT outperformed adapters for adaptation.
@@ 8.2 Sparse Models for NLP
Sparse language models (Evci et al., 2020; Mostafa and Wang, 2019; Dettmers and Zettlemoyer, 2019) route inputs through a subset of the total model parameters. Our work builds most directly on the Branch-Train-Merge (Li et al., 2022; Gururangan et al., 2023) algorithm, which results in full-model experts specialized on domains defined by metadata or a learned clustering. This design expands on early Mixture-of-Experts (MoE) models (Jacobs et al., 1991) and on DEMix layers (Gururangan et al., 2022), which routes sequences to per-layer experts based on metadata.
Other MoE models have recently been applied to multilingual settings. Pfeiffer et al. (2022) develop a multilingual expert model with language-specific routing, and Kudugunta et al. (2021) develop a machine translation model with routing determined by the source-target language pair or the target language. Similar to BTM, Jang et al. (2023) trains experts specialized to different tasks, including five machine-translation language pairs.
@@ 9 Conclusion
This work presents an approach to break the curse of multilinguality by extending sparse language modeling to the multilingual setting with x-elm. We find that x-elms achieve better perplexity on every language over standard, dense language models trained with the same compute budget; expert language models can also be easily adapted to new languages without catastrophic forgetting. x-elms present additional benefits over dense models for multilingual modeling, including training efficiency and flexibility. Finally, we show that these language modeling improvements transfer to downstream, in-context learning performance.
While our experiments show that x-elm outperforms dense LMs, we foresee many avenues of future work to further tailor sparse modeling to multilinguality. These include better methods for data allocation—such as clustering methods that leverage cross-lingual signal— and algorithmic improvements to better allocate compute and more effectively ensemble models. By proving the efficacy of sparse language modeling in the multilingual setting, we hope to inspire future work in this vein that fairly models every language while leveraging the potential of cross-lingual learning.
@@
This work focuses on rigorously examining the effect of training x-elms in a limited number of settings, training languages, and data sources; this is both to ensure that we provide comprehensive comparisons with prior approaches to multilingual language modeling and due to computational limitations. Therefore, the proposed method should be further verified in other settings. In particular, in future work, we hope to examine how x-elm performs at scale when using larger experts, more languages, and larger training budgets. Additionally, while we consider one seed model, XGLM Lin et al. (2022), future work should examine the effect of other pretrained initializations as well as training our own seed models to test how much multilinguality is needed in x-elm initialization.
@@ Appendix A Additional Experimental Details
A.1 Pretraining
A.2 In-Context Learning
We reimplement the evaluation protocol from Lin et al. (2022), where the model scores multiple versions of every example (with the different possible labels filled in), and the label of the highest-scoring version is considered as the model’s prediction. We use the English prompt formats and evaluation protocols developed for the seed LM of our experts, XGLM, for the downstream tasks of XNLI, XStoryCloze, and PAWS-X. The prompt templates we use are reprorted in Table 6.
In the few-shot setting, we perform five evaluation runs with different demonstration samples and reported the average performance. All few-shot experiments are performed with eight random demonstrations. Unless otherwise stated, we evaluate performance on the development set and sample demonstrations from the training set. As we are testing x-elm’s cross-lingual abilities, the demonstrations are in English for every target language.
A.3 Licensing and Intended Use
All of the artifacts used to build the x-elms presented in this work were released for use within academic research Gururangan et al. (2023); Lin et al. (2022); Xue et al. (2021). This also holds for the evaluation benchmarks used to validate the x-elms Xue et al. (2021); Conneau et al. (2018); Lin et al. (2022); Yang et al. (2019). Therefore, the intended use of the code and models presented in this work is in open-source research as well; paricularly, to improve the performance and useablilty of multilingual models for all languages.
Language mC4 Size (%) XGLM Size
@@ A.3 Licensing and Intended Use
Table 5: The frequencies and relative percentages of different languages in our training corpus (an mC4 subsample) and in the XGLM corpus, CC100-XL Lin et al. (2022). Sizes are in gigabytes (GiB). EN, ES, FR, and RU are downsampled to 1,024 shards for mC4.
@@
Appendix B Additional x-elm Analysis
B.1 Comparing the Data Distribution of Clustering Techniques
B.2 Sparse TF-IDF Ensembling
B.3 x-elm Forgetting
We evaluate x-elms as a set of models by dynamically choosing the best expert for a given evaluation setting or ensembling the experts’ outputs. However, each expert is initialized with a model trained on all the languages we consider. This prompts the question: how much do individual experts forget [^12] We consider an expert to have forgetten information about a language if its perplexity on that language increases. about the languages they are not specialized to?
Figure 6: Heatmap comparing individual x-elm perplexities to the seed LM with TF-IDF (left) and Typ. (right). Rows give results for $k=8,4,16$, respectively. Positive scores indicate that the expert forgot that language. For Typ. clusters, languages that the model was explicitly trained on are grayed out.
Forgetting occurs as x-elms become more specialized.
We compare the perplexity of each expert model on all pretraining languages to that of the seed model, XGLM-1.7B (Figure 6).Across the considered values of $k$, we see less forgetting in the x-elms trained on TF-IDF clusters than in those clustered typologically. For the $k=8$ expert setting, the TF-IDF experts only forget on 47.7% of settings, and when forgetting occurs, the perplexity increase over the baseline is 3.10 on average. For typologically clustered experts, these measures are 83.6% and 3.14, respectively.
We observe similar trends for the $k=4$ and $k=16$ x-elms. On average, the $k=4$ TF-IDF experts experience forgetting in only 18.8% of cases with an average perplexity increase of 1.24 when forgetting occurs; the typology experts forget 78.1% of the time with an average perplexity increase of 1.34. For the $k=16$ setting, these statistics are 60.9% and 0.9 for the TF-IDF clusters and 89.4% and 1.24 for the typology clusters.
Lang. k=4 Experts k=16 Experts
XGLM Dense TF-IDFtop1 TF-IDFens Typ. TF-IDFtop1 TF-IDFens Typ.
AR 16.85 15.29 14.99 15.03 15.00 15.60 15.67 15.40
BG 11.31 10.44 10.39 10.39 10.42 11.10 10.70 10.31
DE 15.53 14.02 13.85 13.89 13.71 14.71 14.43 14.5
EL 10.44 9.40 9.36 9.33 9.28 9.72 9.64 9.41
EN 14.37 12.88 12.64 12.71 12.78 13.60 13.23 13.27
ES 16.02 14.13 13.93 13.96 14.06 14.83 14.58 14.59
FR 13.12 11.78 11.62 11.65 11.55 12.38 12.13 12.15
HI 18.28 14.28 14.22 14.21 12.64 16.11 15.67 13.86
JA 14.57 12.31 12.23 12.12 11.73 13.39 13.14 13.18
KO 8.82 7.78 7.81 7.77 7.70 8.14 8.09 7.75
RU 13.43 12.52 12.30 12.33 12.46 12.96 12.76 12.82
SW 19.85 18.70 18.61 18.62 18.19 19.38 19.13 16.43
TR 17.81 15.34 14.85 14.96 14.81 15.67 15.78 15.52
UR 14.38 13.45 13.56 13.73 13.18 13.88 13.87 12.65
VI 13.07 11.39 11.43 11.21 10.32 11.85 11.65 11.59
ZH 17.91 13.74 13.38 13.70 13.11 14.65 14.95 13.58
Avg. 14.74 12.97 12.82 12.85 12.56 13.62 13.46 12.94
Table 7: Per-language and average perplexity results for the $k=4$ and $k=16$ x-elm experiments (original XGLM and $k=1$ dense model included for comparison). Lower numbers are better. Each x-elm setting is trained on 10.5B tokens. TF-IDF ensemble uses more parameters for inference than other evaluations.
x-elms are more likely to forget certain languages.
Appendix C Full Experimental Results
Downstream Evaluation on Individual Languages